OpenStudy (anonymous):

The kings are separated from a deck of standard playing cards and shuffled. One king is randomly selected. What are the odds in favor of drawing a black card?

OpenStudy (anonymous):

There are 4 kings in a deck of cards and separated from the rest. There are 2 kings in black card out of 4 kings separated... so what is the probability of getting one of the black cards in 4 kings...\[p(black ~card)=\frac{2~blacks}{4~cards~of~king}=\frac{1}{2}\]

OpenStudy (anonymous):

thank you. I got the same answer, but my professor counted it wrong. I thought I had misread the question.

OpenStudy (anonymous):

i think we have to consider of separating the 4 kings from the 52 cards and that is \[p(4~kings)=\frac{4}{52}=\frac{1}{13}\]then the odd of getting a king in black suite will be \[p(4~kings~and~in~black~suite)=\frac{1}{2}\times\frac{1}{13}=\frac{1}{26}\]